Find the values of the trigonometric functions of θ from the information given.

Tanθ=-3/4, Cosθ>0

Answer by greenestamps(13198)



Tangent (y/x) is negative in quadrants II and IV; cosine (x/r) is positive in quadrants I and IV. So this angle is in quadrant IV.

With tangent (y/x) equal to -3/4, we can let y=-3 and x=4; the Pythagorean Theorem then gives r=5. Then

sin(t) = y/r = -3/5
cos(t) = x/r = 4/5
tan(t) = y/x = -3/4
cot(t) = 1/tan(t) = -4/3
csc(t) = 1/sin(t) = -5/3
sec(t) = 1/cos(t) = 5/4